OpenStudy (anonymous):

how do we integrate limit 0-2 [20(1-e^-t) dt

OpenStudy (anonymous):

is that a limit? should not be an integral sign?

OpenStudy (anonymous):

\[20\int\limits_{0}^{2}[1 - e^{-t}]dt\] = \[20\int\limits_{0}^{2}dt \] - \[20\int\limits_{0}^{2}e^{-t}dt\]

OpenStudy (anonymous):

20[t]-20[e^-t]/-1 =20[1-0] + 20[e^-1 -1]=20 + 20e^1 -20=20e^-1=20/e